# Noise signatures of a charged Sachdev-Ye-Kitaev dot in mesoscopic transport

### v1

@@ Appendix B Current and noise
$I_{c/h}=\frac{1}{v_{F}}\int d\varepsilon\varepsilon^{n}\rho_{SYK}(\varepsilon)\left[n_{F}(\varepsilon-V_{L},T_{L})-n_{SYK}(\varepsilon-V_{D},T_{D})\right],$ (66)
$S_{c/m/h}=\frac{1}{v_{F}}\int d\varepsilon\varepsilon^{l}\rho_{SYK}(\varepsilon)\left[n_{F}(\varepsilon-V_{L},T_{L})+n_{SYK}(\varepsilon-V_{D},T_{D})-2n_{F}(\varepsilon-V_{L},T_{L})n_{SYK}(\varepsilon-V_{D},T_{D})\right],$ (67)
@@
$\delta S_{l}^{SN}=\left.\frac{\partial S_{l}(\Delta V,\Delta T)}{\partial\Delta V}\right|_{\Delta T=0}=\frac{1}{4Tv_{F}}\int d\varepsilon\rho_{SYK}(\varepsilon)\frac{\varepsilon^{l}\sinh\frac{\varepsilon}{2T}}{\cosh^{3}\frac{\varepsilon}{2T}}.$ (73)
@@
$S_{m}^{JN}=2T\mathcal{L}_{1},\,\,\,\delta S_{m}^{SN}=\mathcal{N}_{1},\,\,\,\delta S_{m}^{\Delta T}=\frac{1}{T}\mathcal{N}_{2};$ (78)
$S_{h}^{JN}=2T\mathcal{L}_{2},\,\,\,\delta S_{h}^{SN}=\mathcal{N}_{2},\,\,\,\delta S_{h}^{\Delta T}=\frac{1}{T}\mathcal{N}_{3}.$ (79)
@@
$\delta S_{h}^{SN}=-\frac{2\textit{i}\pi^{2}T^{3}}{v_{F}}\int_{-\infty}^{\infty}dt\frac{t}{\cosh^{3}\left(\pi Tt\right)}\mathcal{T}\left(\frac{1}{2T}+\textit{i}t\right)-\pi^{2}T^{2}\,\delta S_{c}^{SN}+4T^{2}G_{T},$ (89)
@@
$\delta S_{m}^{SN}=T\,\delta S_{c}^{\Delta T},$ (92)
$\delta S_{m}^{\Delta T}=\frac{1}{T}\,\delta S_{h}^{SN}.$ (93)

### v2

@@ Appendix B Current and noise
$I_{c/h}=\frac{1}{v_{F}}\int d\varepsilon\left(\varepsilon-V_{L}\right)^{n}\rho_{SYK}(\varepsilon)\left[n_{F}(\varepsilon-V_{L},T_{L})-n_{SYK}(\varepsilon-V_{D},T_{D})\right],$ (66)
$S_{c/m/h}=\frac{1}{v_{F}}\int d\varepsilon\left(\varepsilon-V_{L}\right)^{l}\rho_{SYK}(\varepsilon)\left[n_{F}(\varepsilon-V_{L},T_{L})+n_{SYK}(\varepsilon-V_{D},T_{D})-2n_{F}(\varepsilon-V_{L},T_{L})n_{SYK}(\varepsilon-V_{D},T_{D})\right],$ (67)
@@
$\delta S_{l}^{SN}=\left.\frac{\partial S_{l}(\Delta V,\Delta T)}{\partial\Delta V}\right|_{\Delta T=0}=\frac{1}{4Tv_{F}}\int d\varepsilon\rho_{SYK}(\varepsilon)\frac{\varepsilon^{l}\sinh\frac{\varepsilon}{2T}}{\cosh^{3}\frac{\varepsilon}{2T}}-\frac{l}{2v_{k}}\int d\varepsilon\rho_{SYK}(\varepsilon)\frac{\varepsilon^{|l-1|}}{\cosh^{2}\frac{\varepsilon}{2T}}.$ (73)
@@
$S_{m}^{JN}=2T\mathcal{L}_{1},\,\,\,\delta S_{m}^{SN}=\mathcal{N}_{1}-S_{c}^{JN},\,\,\,\delta S_{m}^{\Delta T}=\frac{1}{T}\mathcal{N}_{2};$ (78)
$S_{h}^{JN}=2T\mathcal{L}_{2},\,\,\,\delta S_{h}^{SN}=\mathcal{N}_{2}-2S_{m}^{JN},\,\,\,\delta S_{h}^{\Delta T}=\frac{1}{T}\mathcal{N}_{3}.$ (79)
@@
$\delta S_{h}^{SN}=-\frac{2\textit{i}\pi^{2}T^{3}}{v_{F}}\int_{-\infty}^{\infty}dt\frac{t}{\cosh^{3}\left(\pi Tt\right)}\mathcal{T}\left(\frac{1}{2T}+\textit{i}t\right)-\pi^{2}T^{2}\,\delta S_{c}^{SN},$ (89)
@@
$\delta S_{m}^{SN}=T\,\delta S_{c}^{\Delta T}-S_{c}^{JN},$ (92)
$\delta S_{h}^{SN}=T\delta S_{m}^{\Delta T}-2S_{m}^{JN}.$ (93)